OpenStudy (anonymous):

Completely factor the expression 48x²y + 72xy² + 27y³

myininaya (myininaya):

\[48y \cdot x^2+72y^2 \cdot x +27y^3\] this is in the form of \[Ax^2+Bx+C\] where A=48y B=72y^2 C=27y^3 we want to find two factors of A*C that have product A*C and have sum B. A*C=48y * 27y^3=48(27)y^4=24(54)y^4=12(108)y^4=36(36)y^4 =36y(36y)

myininaya (myininaya):

so B=36y^2+36y^2 so Bx=36xy^2+36xy^2

myininaya (myininaya):

i made a mistake om A*C A*C=36y^2(36y^2)

myininaya (myininaya):

\[48y \cdot x^2+36xy^2+36xy^2+27y^3\] i wil leave the factor by grouping to you